OpenStudy (anonymous):

find y'' (t^(10)lnt)

myininaya (myininaya):

is y=(t^(10)ln(t)) use product rule to find y' \[y'=[t^{10}]' \cdot \ln(t)+t^{10} \cdot [\ln(t)]'\]